OpenStudy (anonymous):

Simplify: -(-1+5)^2 - 2(-4) please show work

OpenStudy (anonymous):

-(4)^2-2(-4) =-16-2(-4) =-16+8 =-8

OpenStudy (anonymous):

thankls